OpenStudy (anonymous):

What is an equation of the line that passes through the point (−2,5) and is perpendicular to the line whose equation is y = ½ x + 5? A. Y=2x+1 B. Y=-2x+1 C. Y=2X+9 D. -2X-9

OpenStudy (tkhunny):

Point is given: (-2,5) Slope is Given: Perpendicular to 1/2 is -2. Do you know why? Ponit-Slope form solve the problem.

OpenStudy (anonymous):

I can't figure it out. I tried point slope and came up with y=1/2x+4...

OpenStudy (tkhunny):

You cannot get "½". That woudl be parallel. (y-5) = -2(x+2) Now what do you get?

OpenStudy (anonymous):

Y=-2x+1?

OpenStudy (tkhunny):

Watch those horrible words very carefully. Perpendicular / Parallel. Not hte same. Promise?

OpenStudy (anonymous):

Haha promise. Is that the correct answer?

OpenStudy (tkhunny):

Got it. Get the next on on the first try, now that you are the master rememberer of wonderful p-words.

OpenStudy (anonymous):

Hahaha, thanks for the help. :)